May 2009 Nightly Log files

22-23 May 2009 LFBG/IRTC Tech Obs 10 - Lost to weather - Closed-dome testing of TCS B27 under CentOS5 and CentOS4
21-22 May 2009 LFBG/IRTC Tech Obs 09 - Lost to weather
20-21 May 2009 LFBG/IRTC Tech Obs 08 - Lost to weather - Closed-dome testing of TCS B26 under CentOS5
19-20 May 2009 LFBG/IRTC Tech Obs 07 - Lost to weather
18-19 May 2009 LFBG/IRTC Tech Obs 06 - Lost to weather
17-18 May 2009 LFBG/IRTC Tech Obs 05 - Offset tests and guiding statistics
16-17 May 2009 LFBG/IRTC Tech Obs 04 - Mostly lost to clouds
15-16 May 2009 LFBG/IRTC Tech Obs 03 - Offset tests, Preset repeatability
14-15 May 2009 LFBG/IRTC Tech Obs 02 - AGW calib, 20-min guiding tests
13-14 May 2009 LFBG/IRTC Tech Obs 01 - Largely lost to technical problems

12-13 May 2009 DG Technical observing night 8 - AGW transformation and offset data
11-12 May 2009 DG Technical observing night 7 - Active Optics verification - nearly photometric
10-11 May 2009 DG Technical observing night 6 - Active Optics calibrations - photometric
09-10 May 2009 DG Technical observing night 5 - Field aberration data and first transformation data - photometric

08-09 May 2009 BG Technical observing night 12 - Transformation OK, more data - photometric
07-08 May 2009 BG Technical observing night 11 - Test Build 26, Problem with Transformation - photometric

06-07 May 2009 DG Technical observing night 4 - Active Optics working over the field
05-06 May 2009 DG Technical observing night 3 - Active Optics working on the y-axis
04-05 May 2009 DG Technical observing night 2 - Guiding working correctly, second half lost to clouds
03-04 May 2009 DG Technical observing night 1 - First images at DG, issues with guiding

-- DougMiller - 29 Sep 2009
Topic revision: r1 - 29 Sep 2009, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback